An equation that defines y as a function of x is given. Complete parts a and b.
5x-3y=1
a) solve the equation for y in terms of x, and replace y with function notation f(x).
Simplify

Answer by stanbon(75887)

An equation that defines y as a function of x is given.
Complete parts a and b.
5x-3y=1
a) solve the equation for y in terms of x, and replace y with function notation f(x).
---------
y = (5/3)x - 1/3
----------------------------
f(x) = (5/3)x - (1/3)
------------------
Cheers,
Stan H.
----------------